# Learning Loss Landscapes in Preference Optimization

### v1

@@ 5 Related Work
Several works in the literature have shown that it is possible to discover machine learning algorithms that outperform algorithms manually designed by researchers (Oh et al., 2020; Lu et al., 2022; Jackson et al., 2024; Alfano et al., 2024). An approach particularly relevant to our method is DiscoPOP by Lu et al. (2024), which leverages an LLM to discover objective functions for LLM tuning. They consider a different space of objective functions from us, as they replace the log-sigmoid in (5) with a generic loss function, following the framework built by Tang et al. (2024). Additionally, instead of searching over a space of parametrized functions, they ask the LLM to generate loss functions in code space. This distinction suggests that our approaches could be complementary, as the model discovered by DiscoPOP could be paired with our learned mirror map. Lastly, DiscoPOP optimizes its objective function directly on the final task, whereas we adopt a two-stage process—optimizing the loss function on a separate task (MuJoCo) and later transferring it to the LLM setting. This transferability underscores the broader applicability of our approach.
@@
A generalization of DPO alternative to ours is $f$-DPO, developed by Wang et al. (2023), which consists in replacing the $\mathtt{KL}$-divergence in (2) with an $f$-divergence and then apply the same heuristic as DPO to obtain the final objective function. We note that the $\mathtt{KL}$-divergence is the only $f$-divergence to be also a Bregman divergence, and vice-versa. They empirically demonstrate that different $f$-divergences lead to different balances between alignment performance and generation diversity, highlighting the trade-offs inherent to this class of algorithms. Huang et al. (2024) further explore this class of PO algorithm and individuate an $f$-divergence for which $f$-DPO is robust to overoptimization.
6 Conclusion
We have introduced a novel framework for Policy Optimization algorithms, as well as a methodology for the automatic discovery of PO algorithms using evolution strategies. Through a systematic evaluation across diverse settings in MuJoCo environments, we demonstrated that our discovered objective functions consistently match or exceed the performance of existing methods, particularly in noisy and mixed-quality datasets where the ORPO baseline struggles. The introduction of temporally aware objective functions further improved performance, allowing the optimization process to vary and adapt during training. Analysing the landscape of the discovered objectives, we give an intuition that justifies improved performance as well as guidance in the design of new PO algorithms. Our results also indicate that the discovered objectives generalize beyond simple reinforcement learning environments, showing promising performance when transferred to LLMs, thereby confirming the broader applicability of our approach.
Appendix A Proof of Theorem 3.1
We provide here a proof for our main result, i.e. Theorem 3.1. The proof to obtain the DPO objective in (5) follow by taking $\phi=e^{x}$
Theorem A.1 (Theorem 3.1).
@@ 6 Conclusion
Appendix B Further discussion of $\omega$-potentials
@@
Appendix C Hyper-parameters
We give the hyper-parameters we use for training in Tables 4 and 5.

### v2

@@ Appendix B Related Work
Several works in the literature have shown that it is possible to discover machine learning algorithms that outperform algorithms manually designed by researchers (Oh et al., 2020; Lu et al., 2022; Jackson et al., 2024; Alfano et al., 2024). An approach particularly relevant to our method is DiscoPOP by (Lu et al., 2024), which leverages an LLM to discover objective functions for LLM tuning. They consider a different space of objective functions from us, as they replace the log-sigmoid in (5) with a generic loss function, following the framework built by (Tang et al., 2024). Additionally, instead of searching over a space of parametrized functions, they ask the LLM to generate loss functions in code space. This distinction suggests that our approaches could be complementary, as the model discovered by DiscoPOP could be paired with our learned mirror map. Lastly, DiscoPOP optimizes its objective function directly on the final task, whereas we adopt a two-stage process—optimizing the loss function on a separate task (MuJoCo) and later transferring it to the LLM setting. This transferability underscores the broader applicability of our approach.
@@
A generalization of DPO alternative to ours is $f$-DPO, developed by (Wang et al., 2023), which consists in replacing the $\mathtt{KL}$-divergence in (1) with an $f$-divergence and then apply the same heuristic as DPO to obtain the final objective function. We note that the $\mathtt{KL}$-divergence is the only $f$-divergence to be also a Bregman divergence, and vice-versa. They empirically demonstrate that different $f$-divergences lead to different balances between alignment performance and generation diversity, highlighting the trade-offs inherent to this class of algorithms. (Huang et al., 2024) further explore this class of PO algorithm and individuate an $f$-divergence for which $f$-DPO is robust to overoptimization.
Appendix C Proof of Theorem 3.1
We provide here a proof for our main result, i.e. Theorem 3.1. The proof to obtain the DPO objective in (5) follows by taking $\phi=e^{x}$.
Theorem C.1 (Theorem 3.1).
@@
Appendix D Further discussion of $\omega$-potentials
@@
Appendix E Further discussion on Evolution Strategies
Evolution Strategies (ES) represent a powerful, backpropagation-free method for optimizing complex functions, that has been particularly successful in the context of long-horizon, noisy, and bi-level optimization tasks such as RL and meta-RL. ES, and in particular the OpenAI-ES algorithm (Salimans et al., 2017), rely on perturbation-based sampling to estimate gradients without requiring backpropagation through the entire computational graph. This feature makes ES well-suited for tasks with long computational graphs, for instance algorithms with many updates, where, due to memory constraints, traditional gradient-based methods have to resort to gradient truncation, introducing bias (Werbos, 1990; Metz et al., 2022; Liu et al., 2022).
In our setting, we use ES to search for the best $\psi$ and $\phi^{-1}$ within the parametrized class introduced in Section 3.1, so that an agent trained using the objective in (10) achieves the highest value. Denote by $\zeta$ the parameters of $\psi$ and $\phi^{-1}$ and by $\pi^{\zeta}$ the final policy obtained optimizing the objective in (10) when using the parametrized $\psi$ and $\phi^{-1}$. Lastly, let $F(\zeta)$ be the expected cumulative reward (or value) of $\pi^{\zeta}$, i.e. $F(\zeta)=\mathbb{E}_{\tau\sim(\mu,\pi^{\zeta},P)}r(\tau)$. At each iteration, we estimate the gradient $\nabla_{\zeta}F(\zeta)$ as
$\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I_{d})}\left[\frac{\epsilon}{2\sigma}(\widehat{F}(\zeta+\sigma\epsilon)-\widehat{F}(\zeta-\sigma\epsilon))\right],$ (12)
where $\mathcal{N}(0,I_{d})$ is the multivariate normal distribution, $d$ is the number of parameters, $\widehat{F}$ is an estimate of $F$, and $\sigma>0$ is a hyperparameter regulating the variance of the perturbations.
Appendix F Further discussion on Online vs Offline Methods
In the domain of RL and preference optimization, the choice between online and offline algorithms presents a critical trade-off, influencing computational efficiency, data requirements, and generalization capabilities. Online methods, such as PPO, iteratively collect and incorporate new data during training. These inherently support exploration of the environment, enabling the discovery of novel strategies or behaviors that are not captured in pre-existing datasets. However, they need feedback for each generated “trajectory” (or response, in the LLM case), which might be expensive to obtain. Online methods are also more complex and particularly sensitive to hyperparameters, often requiring meticulous tuning for stability and efficiency.
Offline algorithms, such as DPO and its variants, rely entirely on pre-collected datasets. These methods are designed for efficiency and simplicity: they don’t require any additional feedback from users and are therefore particularly effective in scenarios where feedback is delayed or unavailable. However, the reliance on static datasets means offline methods may struggle to generalize beyond the training data, particularly if the distribution shift between the training dataset and test time distribution is significant. Additionally, the performance of the algorithm is closely tied to the quality of the training dataset: noisy, biased, or corrupt datasets can severely degrade performance, as these methods cannot mitigate such issues through exploration or resampling.
In summary, RLHF (i.e., online) is considered the superior approach, particularly when substantial amounts of online labels are accessible. This makes it the industry standard (Xu et al., 2024b). While DPO has been theoretically equated to optimizing using PPO and a reward model trained on an offline dataset, recent empirical research (Tang et al., 2024) has challenged this notion. These studies have demonstrated that online methods, such as PPO, consistently outperform offline methods like DPO. This superiority is attributed to the benefits of on-policy sampling.
While DPO has occasionally outperformed PPO, it’s important to note that several studies (Xu et al., 2024b; Song et al., 2024) have consistently shown PPO’s overall superiority. DPO’s relative strength lies in its simpler training regime, which avoids the complexities associated with reward model inaccuracies. However, DPO’s performance is significantly limited by its sensitivity to distribution shift, especially when the offline preference data lacks diversity (Song et al., 2024). This limitation becomes particularly evident when querying the model with out-of-distribution data, a common challenge for methods relying solely on offline data. To mitigate this issue, DPO-iter (Xu et al., 2024b), which incorporates online data, has been proposed as a potential solution.
Appendix G Further discussion on Meta-Learning Algorithms
Meta-learning, or “learning to learn”, has been extensively employed to automate the design of algorithms that can either adapt rapidly with minimal data samples or generalize effectively to unseen data, tasks, or environments. The development of broadly applicable algorithms is particularly critical in the context of preference optimization for LLMs. Here, LLMs are fine-tuned on relatively small datasets of offline data but must generalize to a virtually infinite range of potential user queries. Prior work in meta-learning has demonstrated success in developing generalizable optimization algorithms and loss functions (Lu et al., 2022; Jackson et al., 2024; Lu et al., 2024; Goldie et al., 2024; Kirsch et al., 2020).
At its core, meta-learning is defined as a bilevel optimization problem with an inner and an outer loop. The inner loop consists in an iterative optimization algorithm that trains agents to solve a predetermined task given a set of meta-parameters. The outer loop consists in evaluating the agents trained in the inner loop and update the meta-parameters accordingly, following some optimization method like second order gradient descent (Finn et al., 2017). The evaluation of the agents is typically done on a held-out dataset in supervised learning or by sampling trajectories on the environment simulator in RL (Lu et al., 2022; Jackson et al., 2024). In our setting, the inner loop is the offline preference optimization algorithm, while the outer loop is the agent evaluation on the environment (online) and the update of the meta-parameters $\zeta$.
Base Mixed Quality Noisy $\epsilon=0.1$
DPO 1796 $\pm$ 78 458 $\pm$ 82 693 $\pm$ 131
IPO 2049 $\pm$ 13 1606 $\pm$ 91 739 $\pm$ 118
CPO 2078 $\pm$ 12 1078 $\pm$ 35 1813 $\pm$ 33
ORPO 2022 $\pm$ 15 1039 $\pm$ 20 1710 $\pm$ 33
SimPO 2027 $\pm$ 15 1460 $\pm$ 94 1794 $\pm$ 65
Table 4: Hopper. Perfomance of popular MPO algorithms on the Hopper setting. The agent is randomly initialised.
Appendix H MuJoCo Additional Results
Below, we report a simplified version of the discovered loss on the shuffled and noisy $\epsilon=0.3$ datasets:
$\mathcal{L}(\tau_{w},\tau_{l})=0.82\cdot\mathcal{L}_{\mathrm{SFT}}(\tau_{w})$ (13)
$+1.4\cdot(1.22\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))$
$+0.22\cdot(1.22\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))^{2}$
$+0.2\cdot(1.22\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))^{3}$
$\mathcal{L}(\tau_{w},\tau_{l})=0.82\cdot\mathcal{L}_{\mathrm{SFT}}(\tau_{w})$ (14)
$+\max((1.18\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))^{2},$
$0.1\cdot(1.18\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})))))$
We consider a second set of simulations on MuJoCo, based on the Hopper environment. As for the previous sections, our experiments are implemented in JAX (Bradbury et al., 2018) using the brax (Freeman et al., 2021) and evosax (Lange, 2022) libraries. We report our hyper-parameters in Appendix I.
H.1 Hopper Tasks
Differently from the TLA task, we consider a setting where the agent is randomly initialized and needs to learn a policy from scratch. On Hopper, we train an agent with an expected cumulative reward of 2100 (the expert agent) and an agent with an expected cumulative reward of 900 (the bad agent). We generate the preference datasets in the same way we do for TLA, with the exceptions that the number of rows is 5120 and the trajectories are generated by either the expert or the bad agent. We consider the same three variations of the preference datasets used in TLA, where the expert agent corresponds to the target agent, and the bad agent corresponds to the original agent. We include more data compared to the TLA setting as it takes more datapoints for the agent to learn from scratch rather than to adapt to a slightly different objective (like in the TLA case).
H.2 Further MuJoCo Analsys
In addition to the noisy dataset, we also considered a bad judge setting, where the judge would be more likely to swap the label of a pair of trajectories if their ground truth rewards were closer to each other. This is practically implemented as an increase in the temperature of the Bradley-Terry judge. However, we did not notice significantly different results compared to the simple noisy setting, therefore detailed results are not reported.
H.3 Results
Our Hopper results confirm our conclusions in the TLA setting. All algorithms but DPO come close to matching the performance of the expert agent (2100), with CPO being the best. We can see SimPO is the only algorithm that significantly outperforms the bad agent (performance of 900) in the Mixed Quality setting (the $\gamma$ for SimPO was set very high, $\gamma=10$, as a lower value significantly limited performance).
Appendix I MuJoCo Hyper-parameters
We give the hyper-parameters we use for training.
@@ Appendix I MuJoCo Hyper-parameters
Appendix J LLM Hyper-parameters
Parameter Value
Number of epochs 3
Gradient Accumulation Step 32
Batch Size 2
Total Batch Size 64
Learning rate 5e-5
LoRA Yes
LoRA Rank 128
LoRA Alpha 256
Lora Dropout 0.05
Max length 2048
Table 7: Hyper-parameter settings for LLM Training.